(3a-2a^2)-(6a-3a^2)
1
Expert's answer
2012-05-17T07:17:36-0400
(3a-2a^2)-(6a-3a^2)=3a-2a^2-6a+3a^2=a^2-3a=a(a-3)
Here we& used rules for opening braces. If there is "-" sign before the braces we change sign of each term enclosed. If you see "+" before braces the terms go with signs unchanged.